# An Elementary Proof of the Near Optimality of LogSumExp Smoothing

### v1

@@ Abstract
We consider the design of smoothings of the (coordinate-wise) max function in $\mathbb{R}^{d}$ in the infinity norm. The LogSumExp function $f(x)=\ln(\sum^{d}_{i}\exp(x_{i}))$ provides a classical smoothing, differing from the max function in value by at most $\ln(d)$. We provide an elementary construction of a lower bound, establishing that every overestimating smoothing of the max function must differ by at least $\sim 0.8145\ln(d)$. Hence, LogSumExp is optimal up to constant factors. However, in small dimensions, we provide stronger, exactly optimal smoothings attaining our lower bound, showing that the entropy-based LogSumExp approach to smoothing is not exactly optimal.
@@ 1 Introduction
where $\|f-\sigma\|_{\infty}\coloneqq\sup_{x\in\mathbb{R}^{d}}|f(x)-\sigma(x)|$. Generally, we say that $f$ is a $\delta$-smoothing with respect to the infinity norm of the max function if $f$ is convex and $1$-smooth w.r.t. $\|\cdot\|_{\infty}$ with $\|f-\sigma_{\mathtt{max}}\|_{\infty}\leq\delta$. Hence, $f_{\mathtt{LSE}}$ is a $\ln(d)$-smoothing of $\sigma_{\mathtt{max}}$.
@@
for a $\delta$-smoothing $f$ of $\sigma_{\mathtt{max}}$. Since $f$ is 1-smooth with respect to the infinity norm and each individual $g_{i}$ is smooth and Lipschitz, the objective $y\mapsto\frac{\epsilon}{2\delta}f\circ\frac{2\delta}{\epsilon}g\left(y\right)$ above must be $(L+2\delta M^{2}/\epsilon)$-smooth with respect to the Euclidean norm in $\mathbb{R}^{n}$ [4, Proposition 4.1]. Further, it differs from the original objective by at most $\epsilon/2$. As a result, by applying an accelerated algorithm for smooth convex optimization like Nesterov’s method [5] to produce an $\epsilon/2$-minimizer, an $\epsilon$-minimizer to the original problem can be produced with a first-order oracle complexity of at most
@@
$\delta^{\star}(d)\coloneqq\begin{cases}\min_{f}&\|f-\sigma_{\mathtt{max}}\|_{\infty}\\ \mathrm{s.t.}&f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\text{ is convex and $1$-smooth in $\|\cdot\|_{\infty}$}.\end{cases}$
@@
$\delta^{\star}_{\mathtt{over}}(d)\coloneqq\begin{cases}\min_{f}&\|f-\sigma_{\mathtt{max}}\|_{\infty}\\ \mathrm{s.t.}&f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\text{ is convex and $1$-smooth in $\|\cdot\|_{\infty}$}\\ &f(x)\geq\sigma_{\mathtt{max}}(x)\ \forall x\in\mathbb{R}^{d}.\end{cases}$
@@ 1.1 Our Contributions
First, we consider the question of whether $\delta^{\star}(d)$ necessarily scales with $\ln(d)$. We provide two lower bounds establishing this fact. To establish a baseline grounded in the literature, we first give a reduction to a statistical prediction game via online learning, yielding the following lower bound.
@@
This bound leaves a factor of four gap with the LogSumExp upper bound of $0.5\ln(d)$ and relies on nontrivial statistical arguments of [8]. As our main result, we further improve on this, deriving stronger lower bounds directly via an elementary constructive approach relying only on classical inequalities for smooth convex functions and a combinatorial recurrence. At its core, our analysis works inductively by considering elementary inequalities between certain $j_{\ell}$-sparse points $x^{(j_{\ell})}$. For any increasing integer sequence $j_{\ell}$ from $1$ to $d$ of any length $k$, we derive a lower bound. Optimizing over these bounds, we arrive at the following maximal partition sum quantity
$\gamma(d)\coloneqq\begin{cases}\max&\sum_{\ell=1}^{k}\left(1-\frac{j_{\ell-1}}{j_{\ell}}\right)^{2}\\ \mathrm{s.t.}&1=j_{0}<j_{1}<\dots<j_{k-1}<j_{k}=d.\end{cases}$
Our main result shows that this provides a lower bound on the optimal smoothing gap $\delta^{\star}(d)$.
@@ 1.2 Related Works
Prior Lower Bound Theory for Smoothings. A lower bound proof for $\delta^{\star}(d)$ on the order of $\Omega(\ln(d))$ was previously proposed in [7, Theorem 4.4] via limits on Lipschitz mappings satisfying (1.3). However, their approach relies on a game-theoretic argument involving the existence of Nash equilibria. The needed application of Glicksberg’s theorem is problematic due to the non-compactness of the associated game’s strategy space of $\mathbb{R}^{d}$. As a result, the existence of the required equilibrium remains an open technical gap. Our work here circumvents this gap by using entirely constructive, elementary proof techniques.
@@
then a theory for optimal smoothings of general sublinear functions was given by the authors in [10]. In this case, the optimal difference between a smooth convex $f$ and $\sigma_{\mathtt{max}}$ is bounded by a constant for all $d$ (in particular being at most $1/4$, see [10, Section 4.2.3]). As a result, $f_{\mathtt{LSE}}$ is fundamentally suboptimal not just by constants but in a big-O for two-norm smoothing.
@@
In online convex optimization, algorithms often must have regret scale at least with $\Omega(\sqrt{\ln(d)})$. This is proven by a reduction to coin flipping games [8] and attained up to constant factors by approaches using entropic regularization, dual to the LogSumExp function. See the surveys [13, Section 2.5] and [14, Section 7]. Our proof of Proposition 1.1 reduces to this prior art.
@@
2 Proof of Proposition 1.1 – A Statistical Lower Bound Reduction
$\delta^{\star}(d)=\begin{cases}\min&\frac{1}{2}\mathrm{Range}(h)\\ \mathrm{s.t.}&h\colon\Delta_{d}\rightarrow\mathbb{R}\text{ is $1$-strongly convex in the one-norm}\end{cases}$ (2.1)
The claimed equivalence follows by considering the bijection $h=f^{*}$:
This bijection maps between the two problems’ domains as (i) a convex function $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ has $\|f-\sigma_{\mathtt{max}}\|_{\infty}<\infty$ if and only if $f^{*}$ has domain $\Delta_{d}$ and (ii) a function $f$ is $1$-smooth with respect to the infinity norm if and only if $f^{*}$ is $1$-strongly convex with respect to the one-norm by [17, Theorem 5.26]. The equivalence of objective values under this bijection follows by considering the max function’s representation $\sigma_{\mathtt{max}}(x)=\sup_{\lambda\in\Delta_{d}}\langle\lambda,x\rangle$ as $f(x)=\sup_{\lambda\in\Delta_{d}}\langle\lambda,x\rangle-h(\lambda)$ must have
From this, the claimed lower bound follows directly from the standard lower bounds on “regret” in online learning. Cesa-Bianchi et al. [8, Theorem 3.2.3] provide a statistical argument [^3] Cesa-Bianchi et al. [8]’s lower bound is proven by considering the task of predicting a series of $T$ coin flips with expert advice from $d$ experts. Online learning seeks to make predictions with minimal deviation from the performance of the best expert, i.e., “regret”. If the series of coin flips is entirely fair, then no algorithm can outperform on average, while the best of $d$ random experts will statistically outperform the average. establishing that no algorithm can achieve a regret bound lower than $(1+o(1))\sqrt{\ln(d)T/2}$. As an upper bound, Shalev-Shwartz [13, Theorem 2.11] establishes that for any $1$-strongly convex regularizer $h$ on the simplex, the “Follow-the-Regularized-Leader” algorithm has regret at most $\sqrt{2\mathrm{Range}(h)T}$. Combining these upper and lower bounds on regret, it follows that
Then from the dual equivalence (2.1), $\delta^{\star}(d)\geq(1/8+o(1))\ln(d)$. $\Box$
@@
$\delta^{\star}(d)=\hat{\delta}^{\star}(d)\coloneqq\begin{cases}\min_{f}&\|f-\sigma_{\mathtt{max}}\|_{\infty}\\ \mathrm{s.t.}&f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\text{ is convex and $1$-smooth in $\|\cdot\|_{\infty}$}\\ &f(x)=f(Px)\ \forall x\in\mathbb{R}^{d},P\in\Pi_{d}.\end{cases}$
@@
Clearly $\delta^{\star}(d)\leq\hat{\delta}^{\star}(d)$ as it has further constrained the search for an optimal smoothing. We establish the reverse direction by showing that for any feasible $f$ for $\delta^{\star}(d)$, a permutation invariant, convex, $1$-smooth in the infinity norm $\hat{f}$ can be constructed that is no farther from $\sigma_{\mathtt{max}}$. To this end, consider any convex, $1$-smooth function $f$. Define
@@
It is clear that for each $P\in\Pi_{d}$, the map $x\mapsto f(Px)$ is convex and $1$-smooth in the infinity norm. [^4] Applying the chain rule and the simple observation $\left\|P\right\|_{\infty\to\infty}=1$ gives the smoothness result. As a result, $\hat{f}$ is convex and $1$-smooth in the infinity norm. Finally, we establish $\hat{f}$ is no farther from $\sigma_{\mathtt{max}}$ than $f$ is as
@@
Next, we show that the subdifferential $\partial f(x)\coloneqq\{\zeta\mid f(y)\geq f(x)+\langle\zeta,y-x\rangle\ \forall y\}$ of any convex approximator must lie in the simplex $\Delta_{d}\subset\mathbb{R}^{d}$.
@@
Now let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a convex, permutation invariant function that is $1$-smooth in $\|\cdot\|_{\infty}$ and satisfies $\|f-\sigma_{\mathtt{max}}\|_{\infty}<\infty$. For $j=1,\dots,d$, let $x^{(j)}\in\mathbb{R}^{d}$ denote the point $(1/j,\dots,1/j,0,\dots 0)$ with first $j$ entries set equal to $1/j$. For any $\alpha>0$ and $i,j=1\dots,d$, convexity and infinity norm smoothness of $f$ [17, Theorem 5.8] implies
$Q_{i,j}(\alpha)\coloneqq f(\alpha x^{(i)})-f(\alpha x^{(j)})-\left\langle\nabla f(\alpha x^{(j)}),\alpha x^{(i)}-\alpha x^{(j)}\right\rangle-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{1}\geq 0.$ (3.1)
@@
$Q_{i,j}(\alpha)$ $=f(\alpha x^{(i)})-f(\alpha x^{(j)})-\alpha\left\langle\nabla f(\alpha x^{(j)}),\frac{j}{i}x^{(j)}-x^{(j)}\right\rangle-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{1}$
@@
$\leq f(\alpha x^{(i)})-f(\alpha x^{(j)})-\left(\frac{j}{i}-1\right)\left\langle\nabla f(\alpha x^{(j)}),\alpha x^{(j)}\right\rangle-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{1}$
$=f(\alpha x^{(i)})-\sigma_{\mathtt{max}}(\alpha x^{(i)})-\left[f(\alpha x^{(j)})-\sigma_{\mathtt{max}}(\alpha x^{(j)})\right]-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{1}$
@@
$\leq f(\alpha x^{(i)})-\sigma_{\mathtt{max}}(\alpha x^{(i)})-\left[f(\alpha x^{(j)})-\sigma_{\mathtt{max}}(\alpha x^{(j)})\right]-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{1}$ (3.2)
@@
$\leq f(\alpha x^{(d)})-\sigma_{\mathtt{max}}(\alpha x^{(d)})-[f(\alpha x^{(1)})-\sigma_{\mathtt{max}}(\alpha x^{(1)})]-\sum_{\ell=1}^{k}\frac{1}{2}\|\nabla f(\alpha x^{(j_{\ell})})-\nabla f(\alpha x^{(j_{\ell-1})})\|^{2}_{1}$
$\leq 2\|f-\sigma_{\mathtt{max}}\|_{\infty}-\sum_{\ell=1}^{k}\frac{1}{2}\|\nabla f(\alpha x^{(j_{\ell})})-\nabla f(\alpha x^{(j_{\ell-1})})\|^{2}_{1}$
where the first inequality is by (3.1), the second inequality holds because of inequality (3.2) and the fact that the terms $f(\alpha x^{(j_{\ell})})-\sigma_{\mathtt{max}}(\alpha x^{(j_{\ell})})-\left[f(\alpha x_{j_{\ell}-1})-\sigma_{\mathtt{max}}(\alpha x^{(j_{\ell-1})})\right]$ telescope, while the last inequality follows directly from the definition of $\|f-\sigma_{\mathtt{max}}\|_{\infty}$. Rearranging and taking the limit as $\alpha$ tends to infinity gives
$\|f-\sigma_{\mathtt{max}}\|_{\infty}$ $\geq\lim_{\alpha\to\infty}\sum_{\ell=1}^{k}\frac{1}{4}\|\nabla f(\alpha x^{(j_{\ell})})-\nabla f(\alpha x^{(j_{\ell-1})})\|^{2}_{1}.$ (3.3)
@@
$0\leq\lim_{\alpha\to\infty}\frac{1}{j}-\lambda_{j}(\alpha)=\lim_{\alpha\to\infty}\frac{1}{\alpha}\left(\frac{\alpha}{j}-\alpha\lambda_{j}(\alpha)\right)$ $=\lim_{\alpha\to\infty}\frac{1}{\alpha}\left(\sigma_{\mathtt{max}}(\alpha x^{(j)})-\left\langle\nabla f(\alpha x^{(j)}),\alpha x^{(j)}\right\rangle\right)$
@@
$\|f-\sigma_{\mathtt{max}}\|_{\infty}$ $\geq\sum_{\ell=1}^{k}\frac{1}{4}\|\lim_{\alpha\to\infty}\nabla f(\alpha x^{(j_{\ell})})-\lim_{\alpha\to\infty}\nabla f(\alpha x^{(j_{\ell-1})})\|^{2}_{1}$
$=\sum_{\ell=1}^{k}\frac{1}{4}\|x^{(j_{\ell})}-x^{(j_{\ell-1})}\|^{2}_{1}$
$=\sum_{\ell=1}^{k}\frac{1}{4}\left(j_{\ell-1}\left(\frac{1}{j_{\ell-1}}-\frac{1}{j_{\ell}}\right)+(j_{\ell}-j_{\ell-1})\frac{1}{j_{\ell}}\right)^{2}$
$=\sum_{\ell=1}^{k}\left(1-\frac{j_{\ell-1}}{j_{\ell}}\right)^{2}$
@@
$\|f-\sigma_{\mathtt{max}}\|_{\infty}\geq\begin{cases}\max&\sum_{\ell=1}^{k}\left(1-\frac{j_{\ell-1}}{j_{\ell}}\right)^{2}\\ \mathrm{s.t.}&1=j_{0}<j_{1}<\dots<j_{k-1}<j_{k}=d.\end{cases}$
Finally, we prove our asymptotic claim by showing that
@@
Acknowledgments. This work was supported in part by the Air Force Office of Scientific Research under award number FA9550-23-1-0531. Benjamin Grimmer was additionally supported as a fellow of the Alfred P. Sloan Foundation.

### v2

@@ Abstract
We consider the design of smoothings of the (coordinate-wise) max function in $\mathbb{R}^{d}$ in the infinity norm. The LogSumExp function $f(x)=\ln(\sum^{d}_{i}\exp(x_{i}))$ provides a classical smoothing, differing from the max function in value by at most $\ln(d)$. We provide an elementary construction of a lower bound, establishing that every overestimating smoothing of the max function must differ by at least $\sim 0.8145\ln(d)$. Hence, LogSumExp is optimal up to small constant factors. However, in small dimensions, we provide stronger, exactly optimal smoothings attaining our lower bound, showing that the entropy-based LogSumExp approach to smoothing is not exactly optimal.
@@ 1 Introduction
where $\|f-\sigma\|_{\infty}:=\sup_{x\in\mathbb{R}^{d}}|f(x)-\sigma(x)|$. Generally, we say that $f$ is a $\delta$-smoothing with respect to the infinity norm of the max function if $f$ is convex and $1$-smooth w.r.t. $\|\cdot\|_{\infty}$ with $\|f-\sigma_{\mathtt{max}}\|_{\infty}\leq\delta$. Hence, $f_{\mathtt{LSE}}$ is a $\ln(d)$-smoothing of $\sigma_{\mathtt{max}}$.
@@
for a $\delta$-smoothing $f$ of $\sigma_{\mathtt{max}}$. Since $f$ is 1-smooth with respect to the infinity norm and each individual $g_{i}$ is smooth and Lipschitz, the objective $y\mapsto\frac{\epsilon}{2\delta}f\circ\frac{2\delta}{\epsilon}g\left(y\right)$ above must be $(L+2\delta M^{2}/\epsilon)$-smooth with respect to the Euclidean norm in $\mathbb{R}^{n}$ [2, Proposition 4.1]. Further, it differs from the original objective by at most $\epsilon/2$. As a result, by applying an accelerated algorithm for smooth convex optimization like Nesterov’s method [12] to produce an $\epsilon/2$-minimizer, an $\epsilon$-minimizer to the original problem can be produced with a first-order oracle complexity of at most
@@
$\delta^{\star}(d):=\begin{cases}\min_{f}&\|f-\sigma_{\mathtt{max}}\|_{\infty}\\ \mathrm{s.t.}&f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\text{ is convex and $1$-smooth in $\|\cdot\|_{\infty}$}.\end{cases}$
@@
$\delta^{\star}_{\mathtt{over}}(d):=\begin{cases}\min_{f}&\|f-\sigma_{\mathtt{max}}\|_{\infty}\\ \mathrm{s.t.}&f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\text{ is convex and $1$-smooth in $\|\cdot\|_{\infty}$}\\ &f(x)\geq\sigma_{\mathtt{max}}(x)\ \forall x\in\mathbb{R}^{d}.\end{cases}$
@@ 1.1 Our Contributions
A lower bound on $\delta^{\star}(d)$ of $\ln(d)/8$ can be extracted from two related lower bounds in the literature. Epasto et al. [6] provided a lower bound via limits on any Lipschitz mappings satisfying (1.3). Their approach relies on a game-theoretic argument establishing the existence of a Nash equilibrium from an application of Glicksberg’s theorem and a careful compactification approach.
Proposition 1.1 (Theorem 4.4, Epasto et al. [6]).
For any dimension $d=2^{2^{k}}$ for an integer $k\geq 0$, one has
$\delta^{\star}(d)\geq\frac{1}{8}\ln(d).$
Alternatively, via a reduction to a statistical prediction game of [5] in online learning, one can derive an asymptotically matching lower bound for all $d$ (proven in Appendix A for completeness).
@@
These bounds leave a factor of four gap with the LogSumExp upper bound of $0.5\ln(d)$ and rely on nontrivial equilibrium existence and statistical arguments, respectively. As our main result, we further improve on this, deriving stronger lower bounds directly via an elementary constructive approach relying only on classical inequalities for smooth convex functions and a combinatorial recurrence. At its core, our analysis works inductively by considering elementary inequalities between certain $j_{\ell}$-sparse points $x^{(j_{\ell})}$. For any increasing integer sequence $j_{\ell}$ from $1$ to $d$ of any length $k$, we derive a lower bound. Optimizing over these bounds, we arrive at the following maximal partition sum quantity
$\gamma(d):=\begin{cases}\max&\sum_{\ell=1}^{k}\left(1-\frac{j_{\ell-1}}{j_{\ell}}\right)^{2}\\ \mathrm{s.t.}&1=j_{0}<j_{1}<\dots<j_{k-1}<j_{k}=d.\end{cases}$
Our main result shows that this provides a lower bound on the optimal smoothing gap $\delta^{\star}(d)$. Our proof of Theorem 1.1 in fact provides a stronger lower bound, applying to any $p$-norm smoothing.
@@ 1.2 Related Works
Prior Lower Bound Theory for Smoothings. As discussed above, lower bound proofs can be extracted from the literature [5, 6] via the game theoretic or statistical reductions in Propositions 1.1 and 1.2 respectively. Our work here circumvents the need for such heavy machinery by using entirely constructive, elementary proof techniques. Moreover, our tailored, direct approach yields tighter lower bounds, being exact for $d=2,3$ at least.
@@
then a theory for exactly optimal smoothings of general sublinear functions was given by the authors in [15]. In this case, the optimal difference between a smooth convex $f$ and $\sigma_{\mathtt{max}}$ is bounded by a constant for all $d$ (in particular being at most $1/4$, see [15, Section 4.2.3]). As a result, $f_{\mathtt{LSE}}$ is fundamentally suboptimal not just by constants but in a big-O for two-norm smoothing.
@@
In online convex optimization, algorithms often must have regret scale at least with $\Omega(\sqrt{\ln(d)})$. This is proven by a reduction to coin flipping games [5] and attained up to constant factors by approaches using entropic regularization, dual to the LogSumExp function. See the surveys [16, Section 2.5] and [13, Section 7]. Our proof of Proposition 1.2 reduces to this prior art.
@@
We prove the first result of $\delta\geq\gamma$ more generally than the case of $\infty$-norms of interest to this paper, deriving a bound for any $p$-norm smoothing. Letting $q$ denote the dual norm exponent (i.e., $1/p+1/q=1$), a function is $1$-smooth with respect to $\|\cdot\|_{p}$ if $\|\nabla f(x)-\nabla f(y)\|_{q}\leq\|x-y\|_{p}$. Correspondingly, for any $p$, define
$\delta^{\star}_{p}(d):=\begin{cases}\min_{f}&\|f-\sigma_{\mathtt{max}}\|_{\infty}\\ \mathrm{s.t.}&f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\text{ is convex and $1$-smooth in $\|\cdot\|_{p}$}\end{cases}$
and
$\gamma_{p}(d):=\begin{cases}\max&\sum_{\ell=1}^{k}\frac{1}{4}\left(j_{\ell-1}\left(\frac{1}{j_{\ell-1}}-\frac{1}{j_{\ell}}\right)^{q}+(j_{\ell}-j_{\ell-1})\left(\frac{1}{j_{\ell}}\right)^{q}\right)^{2/q}\\ \mathrm{s.t.}&1=j_{0}<j_{1}<\dots<j_{k-1}<j_{k}=d.\end{cases}$
@@
$\delta^{\star}_{p}(d)=\hat{\delta}^{\star}_{p}(d):=\begin{cases}\min_{f}&\|f-\sigma_{\mathtt{max}}\|_{\infty}\\ \mathrm{s.t.}&f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\text{ is convex and $1$-smooth in $\|\cdot\|_{p}$}\\ &f(x)=f(Px)\ \forall x\in\mathbb{R}^{d},P\in\Pi_{d}.\end{cases}$
@@
Clearly $\delta^{\star}_{p}(d)\leq\hat{\delta}^{\star}_{p}(d)$ as it has further constrained the search for an optimal smoothing. We establish the reverse direction by showing that for any feasible $f$ for $\delta^{\star}(d)$, a permutation invariant, convex, $1$-smooth in the $p$-norm $\hat{f}$ can be constructed that is no farther from $\sigma_{\mathtt{max}}$. To this end, consider any convex, $1$-smooth function $f$ with respect to $\|\cdot\|_{p}$. Define
@@
It is clear that for each $P\in\Pi_{d}$, the map $x\mapsto f(Px)$ is convex and $1$-smooth in the $p$-norm. [^3] Applying the chain rule and the simple observation $\left\|P\right\|_{p\to p}=1$ gives the smoothness result. As a result, $\hat{f}$ is convex and $1$-smooth in the $p$-norm. Finally, we establish $\hat{f}$ is no farther from $\sigma_{\mathtt{max}}$ than $f$ is as
@@
Next, we show that the subdifferential $\partial f(x):=\{\zeta\mid f(y)\geq f(x)+\langle\zeta,y-x\rangle\ \forall y\}$ of any convex approximator must lie in the simplex $\Delta_{d}\subset\mathbb{R}^{d}$.
@@
Now let $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a convex, permutation invariant function that is $1$-smooth in $\|\cdot\|_{p}$ and satisfies $\|f-\sigma_{\mathtt{max}}\|_{\infty}<\infty$. For $j=1,\dots,d$, let $x^{(j)}\in\mathbb{R}^{d}$ denote the point $(1/j,\dots,1/j,0,\dots 0)$ with first $j$ entries set equal to $1/j$. For any $\alpha>0$ and $i,j=1\dots,d$, convexity and $p$-norm smoothness of $f$ [3, Theorem 5.8] implies
$Q_{i,j}(\alpha):=f(\alpha x^{(i)})-f(\alpha x^{(j)})-\left\langle\nabla f(\alpha x^{(j)}),\alpha x^{(i)}-\alpha x^{(j)}\right\rangle-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{q}\geq 0.$ (2.1)
@@
$Q_{i,j}(\alpha)$ $=f(\alpha x^{(i)})-f(\alpha x^{(j)})-\alpha\left\langle\nabla f(\alpha x^{(j)}),\frac{j}{i}x^{(j)}-x^{(j)}\right\rangle-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{q}$
@@
$\leq f(\alpha x^{(i)})-f(\alpha x^{(j)})-\left(\frac{j}{i}-1\right)\left\langle\nabla f(\alpha x^{(j)}),\alpha x^{(j)}\right\rangle-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{q}$
$=f(\alpha x^{(i)})-\sigma_{\mathtt{max}}(\alpha x^{(i)})-\left[f(\alpha x^{(j)})-\sigma_{\mathtt{max}}(\alpha x^{(j)})\right]-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{q}$
@@
$\leq f(\alpha x^{(i)})-\sigma_{\mathtt{max}}(\alpha x^{(i)})-\left[f(\alpha x^{(j)})-\sigma_{\mathtt{max}}(\alpha x^{(j)})\right]-\frac{1}{2}\|\nabla f(\alpha x^{(i)})-\nabla f(\alpha x^{(j)})\|^{2}_{q}$ (2.2)
@@
$\leq f(\alpha x^{(d)})-\sigma_{\mathtt{max}}(\alpha x^{(d)})-[f(\alpha x^{(1)})-\sigma_{\mathtt{max}}(\alpha x^{(1)})]-\sum_{\ell=1}^{k}\frac{1}{2}\|\nabla f(\alpha x^{(j_{\ell})})-\nabla f(\alpha x^{(j_{\ell-1})})\|^{2}_{q}$
$\leq 2\|f-\sigma_{\mathtt{max}}\|_{\infty}-\sum_{\ell=1}^{k}\frac{1}{2}\|\nabla f(\alpha x^{(j_{\ell})})-\nabla f(\alpha x^{(j_{\ell-1})})\|^{2}_{q}$
where the first inequality is by (2.1), the second inequality holds because of inequality (2.2) and the fact that the terms $f(\alpha x^{(j_{\ell})})-\sigma_{\mathtt{max}}(\alpha x^{(j_{\ell})})-\left[f(\alpha x^{(j_{\ell-1})})-\sigma_{\mathtt{max}}(\alpha x^{(j_{\ell-1})})\right]$ telescope, while the last inequality follows directly from the definition of $\|f-\sigma_{\mathtt{max}}\|_{\infty}$. Rearranging and taking the limit as $\alpha$ tends to infinity gives
$\|f-\sigma_{\mathtt{max}}\|_{\infty}$ $\geq\lim_{\alpha\to\infty}\sum_{\ell=1}^{k}\frac{1}{4}\|\nabla f(\alpha x^{(j_{\ell})})-\nabla f(\alpha x^{(j_{\ell-1})})\|^{2}_{q}.$ (2.3)
@@
$0\leq\lim_{\alpha\to\infty}\frac{1}{j}-\lambda_{j}(\alpha)$ $=\lim_{\alpha\to\infty}\frac{1}{\alpha}\left(\frac{\alpha}{j}-\alpha\lambda_{j}(\alpha)\right)$
$=\lim_{\alpha\to\infty}\frac{1}{\alpha}\left(\sigma_{\mathtt{max}}(\alpha x^{(j)})-\left\langle\nabla f(\alpha x^{(j)}),\alpha x^{(j)}\right\rangle\right)$
@@
$\|f-\sigma_{\mathtt{max}}\|_{\infty}$ $\geq\sum_{\ell=1}^{k}\frac{1}{4}\|\lim_{\alpha\to\infty}\nabla f(\alpha x^{(j_{\ell})})-\lim_{\alpha\to\infty}\nabla f(\alpha x^{(j_{\ell-1})})\|^{2}_{q}$
$=\sum_{\ell=1}^{k}\frac{1}{4}\|x^{(j_{\ell})}-x^{(j_{\ell-1})}\|^{2}_{q}$
$=\sum_{\ell=1}^{k}\frac{1}{4}\left(j_{\ell-1}\left(\frac{1}{j_{\ell-1}}-\frac{1}{j_{\ell}}\right)^{q}+(j_{\ell}-j_{\ell-1})\left(\frac{1}{j_{\ell}}\right)^{q}\right)^{2/q}$
@@
$\|f-\sigma_{\mathtt{max}}\|_{\infty}\geq\gamma_{p}(d)=\begin{cases}\max&\sum_{\ell=1}^{k}\frac{1}{4}\left(j_{\ell-1}\left(\frac{1}{j_{\ell-1}}-\frac{1}{j_{\ell}}\right)^{q}+(j_{\ell}-j_{\ell-1})\left(\frac{1}{j_{\ell}}\right)^{q}\right)^{2/q}\\ \mathrm{s.t.}&1=j_{0}<j_{1}<\dots<j_{k-1}<j_{k}=d.\end{cases}$
Hence $\delta^{\star}_{p}\geq\gamma_{p}$ for all $d$. When $p=\infty$ and $q=1$, this objective reduces to $\sum_{\ell=1}^{k}\left(1-\frac{j_{\ell-1}}{j_{\ell}}\right)^{2}$, giving the theorem’s first claim.
Finally, we prove our asymptotic claim for $p=\infty$ by showing that
@@
Acknowledgments.
This work was supported in part by the Air Force Office of Scientific Research under award number FA9550-23-1-0531. Benjamin Grimmer was additionally supported as a fellow of the Alfred P. Sloan Foundation.
Appendix A Proof of Proposition 1.2 – A Statistical Lower Bound Reduction
$\delta^{\star}(d)=\begin{cases}\min&\frac{1}{2}\mathrm{Range}(h)\\ \mathrm{s.t.}&h\colon\Delta_{d}\rightarrow\mathbb{R}\text{ is $1$-strongly convex in the one-norm}\end{cases}$ (A.1)
The claimed equivalence follows by considering the bijection $h=f^{*}$:
This bijection maps between the two problems’ domains as (i) a convex function $f\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ has $\|f-\sigma_{\mathtt{max}}\|_{\infty}<\infty$ if and only if $f^{*}$ has domain $\Delta_{d}$ and (ii) a function $f$ is $1$-smooth with respect to the infinity norm if and only if $f^{*}$ is $1$-strongly convex with respect to the one-norm by [3, Theorem 5.26]. The equivalence of objective values under this bijection follows by considering the max function’s representation $\sigma_{\mathtt{max}}(x)=\sup_{\lambda\in\Delta_{d}}\langle\lambda,x\rangle$ as $f(x)=\sup_{\lambda\in\Delta_{d}}\langle\lambda,x\rangle-h(\lambda)$ must have
From this, the claimed lower bound follows directly from the standard lower bounds on “regret” in online learning. Cesa-Bianchi et al. [5, Theorem 3.2.3] provide a statistical argument [^5] Cesa-Bianchi et al. [5]’s lower bound is proven by considering the task of predicting a series of $T$ coin flips with expert advice from $d$ experts. Online learning seeks to make predictions with minimal deviation from the performance of the best expert, i.e., “regret”. If the series of coin flips is entirely fair, then no algorithm can outperform on average, while the best of $d$ random experts will statistically outperform the average. establishing that no algorithm can achieve a regret bound lower than $(1+o(1))\sqrt{\ln(d)T/2}$. As an upper bound, Shalev-Shwartz [16, Theorem 2.11] establishes that for any $1$-strongly convex regularizer $h$ on the simplex, the “Follow-the-Regularized-Leader” algorithm has regret at most $\sqrt{2\mathrm{Range}(h)T}$. Combining these upper and lower bounds on regret, it follows that
Then from the dual equivalence (A.1), $\delta^{\star}(d)\geq(1/8+o(1))\ln(d)$. $\Box$